# Mitigating Label Noise using Prompt-Based Hyperbolic Meta-Learning in Open-Set Domain Generalization

### v1

@@ 1 Introduction
Open-Set Domain Generalization (OSDG) represents a critical generalization problem that combines the dual challenges of domain shift and category shift during the test phase. In this task, models are required to provide accurate predictions for categories encountered during training while assigning low confidence scores to unseen categories from new domains. OSDG is particularly relevant in real-world applications with diverse and dynamic conditions, such as healthcare [35], security [5], and autonomous driving [22], where new domains and categories frequently arise during deployment. Recent advancements in OSDG have often adopted meta-learning frameworks [66, 57], where training involves simulating various cross-domain tasks to enhance generalization. These methods rely on precise a priori knowledge from source domains and known categories to learn discriminative embeddings that enable accurate predictions in new domains and new categories during testing.
The presence of label noise further exacerbates the challenges of OSDG, as it undermines the reliability of the a priori knowledge derived from the training set. Despite its importance, the issue of label noise in OSDG has been largely overlooked by the research community. While extensive efforts have been made to address label noise in general classification tasks, existing methods such as relabeling approaches [75, 80, 38], data pruning techniques [30, 27, 85, 40], and loss-based noise-agnostic methods [72, 48, 79, 74] focus primarily on identifying and correcting noisy labels, pruning noisy datasets, or optimizing selectively based on loss values. However, these approaches do not account for the additional complexity of generalizing to unseen domains and distinguishing unseen categories, which are essential in OSDG.
To address this gap, we for the first time systematically investigate the impact of label noise on open-set domain generalization. In this work, we introduce two novel benchmarks for OSDG under noisy label settings, based on the widely-used PACS [33] and DigitsDG [81] datasets. We evaluate a wide range of well-established methods from both the OSDG and noisy label learning fields as baselines. Our experiments reveal that these baseline approaches face significant limitations in this new setting. Specifically, OSDG methods are highly sensitive to label noise, while noisy label learning approaches struggle to handle the large domain shifts and label shifts inherent in the train-test split of OSDG task.
@@
  • To the best of our knowledge, this is the first work to explore the task of Open-Set Domain Generalization under Noisy Labels (OSDG-NL). We establish two benchmarks for this purpose, leveraging methodologies from both the open-set domain generalization and label noise learning paradigms, using PACS [33] and DigitsDG [81] datasets.
  • To effectively tackle the challenges inherent to OSDG-NL, we propose a novel approach named HyProMeta, which employs hyperbolic categorical prototypes for label noise-aware meta-learning and introduces new category-aware prompt learning to enhance the model’s generalizability to unseen categories and assist the meta-learning.
@@ 2.1 Open-Set Domain Generalization
Open-set domain generalization consists of two interconnected challenges targeting the generalizability of deep learning models, which are the domain generalization challenge and the open-set recognition challenge. Open-set recognition focuses on rejecting instances of unknown categories during inference by assigning them low confidence scores [20, 53] while domain generalization involves training a deep learning model on multiple source domains and enabling it to generalize effectively to unseen domains at test time [60, 64].
Different approaches are adopted to deal with these two different data shifts existing in train and test sets respectively. Open-set recognition is well-explored by the community through the proposal of techniques such as evidential learning [65, 77, 2], logits calibration [53], GAN-based approaches [31], reconstruction-based methods [73, 25], and reciprocal point-based models [7, 8], while domain generalization strategies aim to bridge the domain gap using a variety of techniques, including data augmentation [67, 51, 83, 21, 82, 37, 39], domain-specific normalization [54], domain adversarial training [19], and GAN-based methods [9, 34].
When tackling the aforementioned two challenges at the same time in open-set domain generalization, most of the research works explored how to achieve good generalizability towards different domains and adopted settings where the known categories are not equally distributed in each source domain [18, 58, 4, 7, 42, 76]. Among them, Katsumata et al. [28] presented a metric learning-based approach to create a discriminative embedding space, aiding open-set domain generalization. Domain-augmented meta-learning is proposed by Shu et al. [57] to tackle the open-set domain generalization problem by introducing more diverse data distributions during training time. Bose et al. [4] introduced ODG-Net, which leverages GANs to synthesize data from a union of training domains, improving cross-domain performance. Wang et al. [66] made notable contributions by formalizing new Open-Set Domain Generalization (OSDG) protocols, where all source domains share the same predefined seen category set. Within these new proposed protocols, meta-learning-based approaches [52, 66] show promising performances when dealing with OSDG tasks. In this work, we adopt the same OSDG settings proposed by Wang et al. [66] on the leveraged DigitsDG and PACS datasets. Our proposed new method, HyProMeta, also adopts meta-learning as the basic framework to deal with open-set domain generalization while incorporating the label noise agnostic learning enabled by hyperbolic space prototypes and new-category aware prompt learning to enhance the model’s generalizability.
@@ 2.2 Noisy Label Learning
Optimizing models with accurate labels is crucial for most deep learning methods to enable them to learn precise a priori knowledge from the provided samples [72]. In contrast, incorrect labels can misguide the optimization process [13]. To address the negative impact of label noise, researchers have proposed various approaches from different perspectives. For instance, methods introduced in [71, 61, 87, 88, 41] aim to identify label noise by modeling the probability of label corruption. Other approaches, such as those in [59, 70, 11], focus on detecting label noise prior to training.
@@
Alternatively, sample re-weighting techniques mitigate the influence of mislabeled samples by assigning lower weights to their loss values [46]. Meta-learning-based methods [56, 68, 79] have also shown promise in handling noisy labels by dynamically adapting the loss function and learning optimal training strategies.
@@
Dynamic sample selection approaches adopt a semi-supervised framework to manage noisy data. These methods begin training with a subset of clean samples and incrementally incorporate mislabeled data in a controlled manner, enabling the model to adapt effectively over time [23, 10, 36]. For instance, TCL [26] employs contrastive learning to develop robust representations and uses a Gaussian Mixture Model for label mapping. Similarly, PLM [78] utilizes part-level labels to guide the model in extracting richer and more detailed information.
@@
Several works have also explored noise-robust loss functions [47, 49, 86] to improve optimization in the presence of label noise. Other studies have focused on employing proper regularization techniques to minimize the adverse effects of noisy labels [69, 12, 45]. For example, BadLabel [75] implements a label-flipping attack, creating indistinguishable loss values between clean and noisy labels. Meanwhile, LSL [29] leverages structural label information to improve learning from noisy data.
@@
Despite these advancements, most existing approaches for learning with label noise fail to address the challenges posed by domain shifts and label distribution shifts encountered at test time. This limitation presents significant robustness challenges for handling noisy labels in the field of Open-Set Domain Generalization (OSDG). In this work, we, for the first time, focus on the task of Open-Set Domain Generalization under Noisy Labels (OSDG-NL). We introduce two novel benchmarks based on the well-established OSDG datasets, namely PACS [33] and DigitsDG [81], by incorporating various levels of label noise. Additionally, we select diverse methods from the OSDG and label noise learning fields to serve as baselines. We further propose a novel approach, HyProMeta, which implements label noise-aware meta-learning and demonstrates superior performance compared to existing methods on the challenging OSDG-NL task.
@@ 3.1 Task Description
In our task, various domains $\Omega_{\mathcal{D}}=\{\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{N_{S}}\}$ are available, we follow the leave one out setting in [66] by choosing $\mathcal{D}_{m}$ as the test domain, while the remaining domains from $\Omega_{\mathcal{D}_{S}}=\Omega_{\mathcal{D}}/\{\mathcal{D}_{m}\}$ serve as source domains during test phase. Let $\Omega_{l}$ denote the set of labels existing in a dataset. To achieve open-set recognition, $\Omega_{u}\subset\Omega_{l}$ are chosen as unseen categories during test phase, while $\Omega_{k}\subset\Omega_{l}$ are known during training, where $\Omega_{l}=\Omega_{u}\cup\Omega_{k}$. The sample from each source domain consists of data-label pairs $(\mathbf{x}_{s},\mathbf{y}_{s})$, where $\mathbf{x}_{s}$ denotes the data and $\mathbf{y}_{s}$ denotes the label. In the noisy label setting, the labels of seen categories in the source domains are perturbed and denoted by $\hat{\mathbf{y}}_{s}$. The goal of our task is to ask the model to give accurate classification results on seen categories for $\mathbf{X}_{t}^{k}=\{(\mathbf{x}_{t}^{k},\mathbf{y}_{t}^{k})|\mathbf{y}_{t}^{k}\in\Omega_{k},\mathbf{X}_{t}^{u}\in\mathcal{D}_{m}\}$ while delivering low confidence score on unseen categories for $\mathbf{X}_{t}^{u}=\{(\mathbf{x}_{t}^{u},\mathbf{y}_{t}^{u})|\mathbf{y}_{t}^{u}\in\Omega_{u},\mathbf{x}_{t}^{u}\in\mathcal{D}_{m}\}$ in the test set, based on the a priori knowledge provided by samples with label noise $\{\hat{\mathbf{y}}_{s}\}$ from known categories in the source domains provided by the train set.
@@ 3.2 Noisy Label Settings
To construct the first OSDG-NL benchmark, we select two types of label noise, which are symmetric and asymmetric label noises. Symmetric noise involves randomly reassigning training labels to other classes at predefined noise levels ($20\%$, $50\%$, $80\%$ in our work), simulating varying degrees of label corruption. In contrast, asymmetric noise considers the semantic similarity between classes, altering labels more likely to the most related categories (e.g., “horse” might be mislabeled as “dog”). We employ the BERT model [15] to extract semantic features based on the textual description of each class and measure the semantic categorical distance using cosine similarity. Then, we use the calculated similarity scores as weights for mislabeling. For asymmetric noise, we set the label noise ratio as $50\%$.
@@ 3.3 Datasets
PACS: The PACS [33] is a well-established dataset in the OSDG field, which consists of $4$ distinct domains: Photo, Art Painting, Cartoon, and Sketch. These domains introduce large visual domain shifts, making PACS suitable for our task. The dataset includes $9,991$ images across $7$ shared categories (i.e., dog, elephant, giraffe, guitar, house, horse, person). PACS is widely used in domain generalization studies with a leave-one-domain-out protocol, which is also leveraged in our work, where the model is trained on three domains and tested on the other domain. We follow the open-set protocol introduced by Wang et al. [66] by selecting the last $1$ category as the unseen category, where all the categories are rearranged into alphabet order according to their text descriptions.
DigitsDG: The Digits-DG [81] is another domain generalization dataset, comprising $4$ distinct domains, i.e., mnist, mnistm, svhn, and syn. Each domain introduces distinct visual styles, ranging from grayscale handwritten digits to real-world and synthetic digit images, resulting in obvious domain shifts. The dataset includes $10$ shared classes (digits $0{\sim}9$). We follow the open-set protocol introduced by Wang et al. [66] by selecting the last $4$ categories as unseen categories which are rearranged into alphabet order.
@@ 3.4 Baselines
Open-set domain generalization baselines: We select $6$ baselines from the open-set domain generalization field, each exemplifying distinct strategies to address domain generalization challenges. These methods cover a wide spectrum of approaches, including meta-learning (MEDIC [66] and MLDG [56]), adversarial learning (ARPL [7]), data augmentation (MixStyle [83]), generative modeling (ODGNet [4]), and optimization techniques (SWAD [6]), ensuring a comprehensive evaluation. These baselines collectively provide a well-rounded comparison, covering key contributions of open-set domain generalization. Thereby we would like to evaluate their performances when they are facing various label noise challenges.
Noisy label learning baselines: We select $6$ recent representative approaches from the noisy label learning field on the general image classification task as another group of baselines in our benchmarks, including TCL [26], NPN [55], BadLabel [75], DISC [43], LSL [29], and PLM [78], each addressing noisy labels through distinct strategies, e.g., sample selection-based contrastive learning, loss-based label denoising, and label correction. These methods provide a comprehensive basis for evaluating noisy label learning techniques.
Figure 2: An overview of the proposed HyProMeta framework. HyProMeta is established on domain-agnostic category-wise prototypes in the hyperbolic space to achieve meta-task assignments. Prompt-based mixed categorical augmentation is leveraged to improve the generalizability of the model toward new data distribution.

### v2

@@ 1 Introduction
Open-Set Domain Generalization (OSDG) represents a critical generalization problem that combines the dual challenges of domain shift and category shift during the test phase. In this task, models are required to provide accurate predictions for categories encountered during training while assigning low confidence scores to unseen categories from new domains. OSDG is particularly relevant in real-world applications with diverse and dynamic conditions, such as healthcare [li2020domain], security [busto2018open], and autonomous driving [guo2022simt], where new domains and categories frequently arise during deployment. Recent advancements in OSDG have often adopted meta-learning frameworks [wang2023generalizable, shu2021open], where training involves simulating various cross-domain tasks to enhance generalization. These methods rely on precise a priori knowledge from source domains and known categories to learn discriminative embeddings that enable accurate predictions in new domains and new categories during testing.
The presence of label noise further exacerbates the challenges of OSDG, as it undermines the reliability of the a priori knowledge derived from the training set. Despite its importance, the issue of label noise in OSDG has been largely overlooked by the research community. While extensive efforts have been made to address label noise in general classification tasks, existing methods such as relabeling approaches [zhang2024badlabel, zheng2020error, li2024nicest], data pruning techniques [kim2021fine, karim2022unicon, zhou2020robust, li2022selective], and loss-based noise-agnostic methods [xu2024skeleton, lukasik2020does, zheng2021meta, yue2024ctrl] focus primarily on identifying and correcting noisy labels, pruning noisy datasets, or optimizing selectively based on loss values. However, these approaches do not account for the additional complexity of generalizing to unseen domains and distinguishing unseen categories, which are essential in OSDG.
To address this gap, we for the first time systematically investigate the impact of label noise on open-set domain generalization. In this work, we introduce two novel benchmarks for OSDG under noisy label settings, based on the widely-used PACS [li2017deeper] and DigitsDG [zhou2020deep] datasets. We evaluate a wide range of well-established methods from both the OSDG and noisy label learning fields as baselines. Our experiments reveal that these baseline approaches face significant limitations in this new setting. Specifically, OSDG methods are highly sensitive to label noise, while noisy label learning approaches struggle to handle the large domain shifts and label shifts inherent in the train-test split of the OSDG task.
@@
  • To the best of our knowledge, this is the first work to explore the task of Open-Set Domain Generalization under Noisy Labels (OSDG-NL). We establish two benchmarks for this purpose, leveraging methodologies from both the open-set domain generalization and label noise learning paradigms, using PACS [li2017deeper] and DigitsDG [zhou2020deep] datasets.
  • To effectively tackle the challenges inherent to OSDG-NL, we propose a novel approach named HyProMeta, which employs hyperbolic category prototypes for label noise-aware meta-learning and introduces new category-aware prompt learning to enhance the model’s generalizability to unseen categories and assist the meta-learning.
@@ 2.1 Open-Set Domain Generalization
Open-set domain generalization consists of two interconnected challenges targeting the generalizability of deep learning models, which are the domain generalization challenge and the open-set recognition challenge. Open-set recognition focuses on rejecting instances of unknown categories during inference by assigning them low confidence scores [geng2020recent, peng2024navigating] while domain generalization involves training a deep learning model on multiple source domains and enabling it to generalize effectively to unseen domains at test time [sun2016deep, wang2022generalizing].
Different approaches are adopted to deal with these two different data shifts existing in the training and testing sets respectively. Open-set recognition is well-explored by the community through the proposal of techniques such as evidential learning [wang2024towards, zhao2023open, bao2021evidential], logits calibration [peng2024navigating], GAN-based approaches [kong2021opengan], reconstruction-based methods [yoshihashi2019classification, huang2022class], and reciprocal point-based models [chen2021adversarial, chen2020learning], while domain generalization strategies aim to bridge the domain gap using a variety of techniques, including data augmentation [wang2020heterogeneous, nam2021reducing, zhou2020domain, guo2023aloft, zhou2020learning, li2021progressive, li2021simple], domain-specific normalization [seo2020learning], domain adversarial training [ganin2016domain], and GAN-based methods [chen2022discriminative, li2024vocoder].
When tackling the aforementioned two challenges at the same time in open-set domain generalization, most of the research works explored how to achieve good generalizability towards different domains and adopted settings where the known categories are not equally distributed in each source domain [fu2020learning, singha2024unknown, bose2023beyond, chen2021adversarial, li2018deep, zhao2022adaptive]. Among them, Katsumata et al. [katsumata2021open] presented a metric learning-based approach to create a discriminative embedding space, aiding open-set domain generalization. Domain-augmented meta-learning is proposed by Shu et al. [shu2021open] to tackle the open-set domain generalization problem by introducing more diverse data distributions during training time. Bose et al. [bose2023beyond] introduced ODG-Net, which leverages GANs to synthesize data from a union of training domains, improving cross-domain performance. Wang et al. [wang2023generalizable] made notable contributions by formalizing new Open-Set Domain Generalization (OSDG) protocols, where all source domains share the same predefined seen category set. Within these new proposed protocols, meta-learning-based approaches [peng2024advancing, wang2023generalizable] show promising performances when dealing with OSDG tasks. In this work, we adopt the same OSDG settings proposed by Wang et al. [wang2023generalizable] on the leveraged DigitsDG and PACS datasets. Our proposed new method, HyProMeta, also adopts meta-learning as the basic framework to deal with open-set domain generalization while incorporating the label noise agnostic learning enabled by hyperbolic space prototypes and new-category aware prompt learning to enhance the model’s generalizability.
@@ 2.2 Noisy Label Learning
Optimizing models with accurate labels is crucial for most deep learning methods to enable them to learn precise a priori knowledge from the provided samples [xu2024skeleton]. In contrast, incorrect labels can misguide the optimization process [cheng2020weakly]. To address the negative impact of label noise, researchers have proposed various approaches from different perspectives. For instance, methods introduced in [xia2019anchor, tanno2019learning, zhu2021clusterability, zhu2022beyond, li2022estimating] aim to identify label noise by modeling the probability of label corruption. Other approaches, such as those in [song2019selfie, wei2021smooth, chen2021beyond], focus on detecting label noise prior to training.
@@
Alternatively, sample re-weighting techniques mitigate the influence of mislabeled samples by assigning lower weights to their loss values [liu2015classification]. Meta-learning-based methods [shu2019meta, wang2020training, zheng2021meta] have also shown promise in handling noisy labels by dynamically adapting the loss function and learning optimal training strategies.
@@
Dynamic sample selection approaches adopt a semi-supervised framework to manage noisy data. These methods begin training with a subset of clean samples and incrementally incorporate mislabeled data in a controlled manner, enabling the model to adapt effectively over time [Han_Yao_Yu_Niu_Xu_Hu_Tsang_Sugiyama_2017, chen2019understanding, li2020dividemix]. For instance, TCL [huang2023twin] employs contrastive learning to develop robust representations and uses a Gaussian Mixture Model for label mapping. Similarly, PLM [zhao2024estimating] utilizes part-level labels to guide the model in extracting richer and more detailed information.
@@
Several works have also explored noise-robust loss functions [liu2020peer, ma2020normalized, zhu2021second] to improve optimization in the presence of label noise. Other studies have focused on employing proper regularization techniques to minimize the adverse effects of noisy labels [wei2021open, cheng2021mitigating, liu2022robust]. For example, BadLabel [zhang2024badlabel] implements a label-flipping attack, creating indistinguishable loss values between clean and noisy labels. Meanwhile, LSL [kim2024learning] leverages structural label information to improve learning from noisy data.
@@
Despite these advancements, most existing approaches for learning with label noise fail to address the challenges posed by domain shifts and label distribution shifts encountered at test time. This limitation presents significant robustness challenges for handling noisy labels in the field of Open-Set Domain Generalization (OSDG). In this work, we, for the first time, focus on the task of Open-Set Domain Generalization under Noisy Labels (OSDG-NL). We introduce two novel benchmarks based on the well-established OSDG datasets, namely PACS [li2017deeper] and DigitsDG [zhou2020deep], by incorporating various levels of label noise. Additionally, we select diverse methods from the OSDG and label noise learning fields to serve as baselines. We further propose a novel approach, HyProMeta, which implements label noise-aware meta-learning and demonstrates superior performance compared to existing methods on the challenging OSDG-NL task.
@@ 3.1 Task Description
In our task, various domains $\Omega_{\mathcal{D}}=\{\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{N_{S}}\}$ are available, we follow the leave one out setting in [wang2023generalizable] by choosing $\mathcal{D}_{m}$ as the test domain, while the remaining domains from $\Omega_{\mathcal{D}_{S}}=\Omega_{\mathcal{D}}/\{\mathcal{D}_{m}\}$ serve as source domains during test phase. Let $\Omega_{l}$ denote the set of labels existing in a dataset. To achieve open-set recognition, $\Omega_{u}\subset\Omega_{l}$ are chosen as unseen categories during test phase, while $\Omega_{k}\subset\Omega_{l}$ are known during training, where $\Omega_{l}=\Omega_{u}\cup\Omega_{k}$. The sample from each source domain consists of data-label pairs $(\mathbf{x}_{s},\mathbf{y}_{s})$, where $\mathbf{x}_{s}$ denotes the data and $\mathbf{y}_{s}$ denotes the label. In the noisy label setting, the labels of seen categories in the source domains are perturbed and denoted by $\hat{\mathbf{y}}_{s}$. The goal of our task is to ask the model to give accurate classification results on seen categories for $\mathbf{X}_{t}^{k}=\{(\mathbf{x}_{t}^{k},\mathbf{y}_{t}^{k})|\mathbf{y}_{t}^{k}\in\Omega_{k},\mathbf{X}_{t}^{u}\in\mathcal{D}_{m}\}$ while delivering low confidence score on unseen categories for $\mathbf{X}_{t}^{u}=\{(\mathbf{x}_{t}^{u},\mathbf{y}_{t}^{u})|\mathbf{y}_{t}^{u}\in\Omega_{u},\mathbf{x}_{t}^{u}\in\mathcal{D}_{m}\}$ in the test set, based on the a priori knowledge provided by samples with label noise $\{\hat{\mathbf{y}}_{s}\}$ from known categories in the source domains provided by the train set.
@@ 3.2 Noisy Label Settings
To construct the first OSDG-NL benchmark, we select two types of label noise, which are symmetric and asymmetric label noises. Symmetric noise involves randomly reassigning training labels to other classes at predefined noise levels ($20\%$, $50\%$, $80\%$ in our work), simulating varying degrees of label corruption. In contrast, asymmetric noise considers the semantic similarity between classes, altering labels more likely to the most related categories (e.g., “horse” might be mislabeled as “dog”). We employ the BERT model [BERT] to extract semantic features based on the textual description of each class and measure the semantic categorical distance using cosine similarity. Then, we use the calculated similarity scores as weights for mislabeling. For asymmetric noise, we set the label noise ratio as $50\%$.
@@ 3.3 Datasets
PACS: The PACS [li2017deeper] is a well-established dataset in the OSDG field, which consists of $4$ distinct domains: Photo, Art Painting, Cartoon, and Sketch. These domains introduce large visual domain shifts, making PACS suitable for our task. The dataset includes $9,991$ images across $7$ shared categories (i.e., dog, elephant, giraffe, guitar, house, horse, person). PACS is widely used in domain generalization studies with a leave-one-domain-out protocol, which is also leveraged in our work, where the model is trained on three domains and tested on the other domain. We follow the open-set protocol introduced by Wang et al. [wang2023generalizable] by selecting the last $1$ category as the unseen category, where all the categories are rearranged into alphabet order according to their text descriptions.
DigitsDG: The Digits-DG [zhou2020deep] is another domain generalization dataset, comprising $4$ distinct domains, i.e., mnist, mnistm, svhn, and syn. Each domain introduces distinct visual styles, ranging from grayscale handwritten digits to real-world and synthetic digit images, resulting in obvious domain shifts. The dataset includes $10$ shared classes (digits $0{\sim}9$). We follow the open-set protocol introduced by Wang et al. [wang2023generalizable] by selecting the last $4$ categories as unseen categories which are rearranged into alphabet order.
DomainNet: DomainNet [peng2019moment] is a large-scale benchmark dataset designed for domain adaptation and generalization tasks, containing approximately $600,000$ images across six distinct domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch. It includes $345$ object categories shared among all domains, enabling evaluation of model performance under significant domain shifts. The first $182$ categories following alphabet order are selected as seen classes while the rest are selected as unseen classes.
@@ 3.4 Baselines
Open-set domain generalization baselines: We select $6$ baselines from the open-set domain generalization field, each exemplifying distinct strategies to address domain generalization challenges. These methods cover a wide spectrum of approaches, including meta-learning (MEDIC [wang2023generalizable], EBiL-HaDS [peng2024advancing] and MLDG [shu2019meta]), adversarial learning (ARPL [chen2021adversarial]), data augmentation (MixStyle [zhou2020domain]), generative modeling (ODGNet [bose2023beyond]), and optimization techniques (SWAD [cha2021swad]), ensuring a comprehensive evaluation. These baselines collectively provide a well-rounded comparison, covering key contributions of open-set domain generalization. Thereby, we would like to evaluate their performances when they are facing various label noise challenges.
Noisy label learning baselines: We select $6$ recent representative approaches from the noisy label learning field on the general image classification task as another group of baselines in our benchmarks, including TCL [huang2023twin], NPN [sheng2024adaptive], BadLabel [zhang2024badlabel], DISC [li2023disc], LSL [kim2024learning], and PLM [zhao2024estimating], each addressing noisy labels through distinct strategies, e.g., sample selection-based contrastive learning, loss-based label denoising, and label correction. These methods provide a comprehensive basis for evaluating noisy label learning techniques.
Figure 2: Overview of the proposed HyProMeta framework for Open-Set Domain Generalization under Noisy Labels (OSDG-NL). The model first computes hyperbolic category-wise prototypes from source domain data and uses the mode of hyperbolic distances to partition samples into clean and noisy subsets. Clean samples are used directly for meta-training, while noisy samples are relabeled based on proximity to prototypes. A learnable prompt is introduced for category-aware augmentation, where mixed samples from different classes are combined with prompt patches and assigned to an auxiliary out-of-distribution class. The framework iteratively updates through meta-train and meta-test phases, enabling robust generalization to unseen categories and domains despite label noise.